OpenStudy (anonymous):

did we do 2/x + 4/2x = -2 ?

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

did we work on that one?

OpenStudy (anonymous):

idk..well we will do it again :P tell me wat u get and ill say if ur right

OpenStudy (anonymous):

2 ?

OpenStudy (anonymous):

the answer is -2, by chance did u forget the sign?

OpenStudy (anonymous):

oh yeah i did get that !

OpenStudy (anonymous):

4/n - 1/2 = -6/3n is it 4 ?

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

i got n=12

OpenStudy (anonymous):

you probably right lol

OpenStudy (anonymous):

4/n - 1/2 = -6/3n 8-n/2n = -6/3n 8 - n = -4 n = 12 understand?

OpenStudy (anonymous):

but if its 8-n =-4 you have to add the 8 to -4 dont you

OpenStudy (anonymous):

no cause 8-n =-4 (8 - n) -8 = -4 -8 -n = -12

OpenStudy (anonymous):

ohhh wait so -12 or 12

OpenStudy (anonymous):

well -n = -12 so n = 12

OpenStudy (anonymous):

ok now x-4/4 +x/3 =6

OpenStudy (anonymous):

u do it and ill correct it :)